At its best possible output level, a firm has total revenue of Rs. 7000 per day and total cost of Rs. 14000 per day. What should this firm do in the short run if: 05
(Hint: See normal and abnormal loss)
i) The firm has total fixed cost of Rs. 6000 per day?
ii) The firm has total variable cost of Rs. 6000 per day?
1
Expert's answer
2020-06-12T13:53:57-0400

"TR=7000"

"TC=14000"

"P=TR-TC"

"=7000-14000=-7000"

i) The firm realises a loss since total revenue is less than the total cost. It is likely to be the case since the total fixed cost does not change with increase or decrease in output hence a normal output with normal cost of normal output.

Abnormal loss"=" ( Normal cost of normal output) "\/" (Normal output)multiplied by the abnormal loss quantity.

In the short run; Total cost"=" Total fixed cost "+" Total variable cost.

In the long run; Total fixed cost "=0" .

Therefore, Total cost "=" Total variable cost.

Total fixed cost remains the same and has to be incurred even if the level of output is zero.

ii) Total variable cost is the opportunity cost incurred in the short run production by a firm that depends on quantity of output. If more of output is produced, the total variable cost increases."TVC" is guided by the law of diminishing marginal returns. In the short run, as more units of variable inputs are added to fixed amounts, the change in total output raises and eventually falls.

At initial stages of production, there is a scope of efficient utilization of fixed factors by using more variable factors. As the variable input employed increases, the productive efficiency of variable inputs ensures that the Total Variable Cost increases at a diminishing rate.